DM. Let a and b be two cardinal numbers. Modify Cantor’s definition of a < b to define a ≤ b. (Hint: Examine what happens if you drop condition (a) from Cantor’s definition of a < b.) 2. Prove that a ≤ a. 3. Prove that if a ≤ b and b ≤ c, then a ≤ c. 4. Do you think that a ≤ b and b ≤ a imply


a = b? Explain your reasoning. (Hint: This is not as trivial as it might look.)



1
Expert's answer
2022-01-04T16:43:51-0500

1.

If for two aggregates M and N with the cardinal numbers a and b , with the condition:

There is a part N1 of N, such that N1 ∼ M,

is fulfilled, it is obvious that this condition still hold if in them M and N are replaced by two equivalent aggregates M0 and N0 . Thus it express a definite relation of the cardinal numbers a and b to one another.


2.

Since a = a, then a = a or a < a

so, "a\le a"


3.

if a ≤ b, then "b=a+k_1,k_1\ge 0"

if b ≤ c, then "c=b+k_2,k_2\ge 0"

so, "c=a+k_1+k_2\implies a\le c"


4

if a ≤ b, then "b=a+k_1,k_1\ge 0"

if b ≤ a, then "a=b+k_2,k_2\ge 0"

so, "b=b+k_1+k_2\implies k_1+k_2=0"

"\implies k_1=k_2=0\implies a=b"